1)Use the Henderson–Hasselbalch equation to calculate the pH of:

a solution that is 13.5 g of HF and 25.0 g of NaF in 125 mL of solution

Express your answer using two decimal places.

2)

A volume of 20.0 mL of a 0.380 M HNO3 solution is titrated with 0.850 M KOH. Calculate the volume of KOH required to reach the equivalence point.

Express your answer to three significant figures and include the appropriate units.
Answer #1

olution 3-5-3 억 HF and aS.o 3.59 V in L mo M=0,675 5-4 м mo = O. 575 = 4,8 M. O,125 iacrd) 3.2 + log 14-钉 [54] =3-15 a) O ESo